# https://github.com/krulllab/COSDD

### v2

@@ 4 Method
Figure 2: a): A Variational Lossy Autoencoder [4] (solid arrows) is trained to model the distribution of noisy images $\mathbf{x}$. The autoregressive decoder models the noise component of the images while the latent code models only the clean signal component $\mathbf{s}$. In a second step (dashed arrows), our novel signal decoder is trained to map latent variables into image space, producing an estimate of the signal underlying $\mathbf{x}$. b): To ensure that the decoder models only imaging noise and the latent code captures only the signal, we modify decoder’s receptive field. While a full autoregressive decoder includes pixels above and to the left (shown in blue), our decoder’s receptive field corresponds to the row-correlated structure of the imaging noise and includes only pixels in the same row as the pixel being modeled.
We propose a VAE-based unsupervised image denoiser for noise that is both signal-dependent and correlated along rows or columns of pixels. It is trained using only noisy images and does not require a pre-trained noise model. We restrict the receptive field of the decoder so it can model the correlations of the noise content but not the correlations of the underlying signal. Following the insights of Chen et al. [4] (Sec. 2.3), the decoder will therefore learn to model the noise, leaving the underlying signal to be encoded in the latent variable $\mathbf{z}$. We then propose a method for taking these latent variables and mapping them back into image space to obtain denoised estimates of clean signals. A full outline of our method can be found in Fig. 2.
4.1 Autoregressive receptive fields for noise
Our AR decoder has a 1-dimensional receptive field that is sufficient for modeling row/column correlated noise (Eq. 2) as it commonly occurs in EMCCD, sCMOS, microbolometer detectors, SCM or STEM (see Sec. 1) by spanning pixels in the same row or column as $x_{i,j}$. See Fig. 2b for a visual representation.
@@ 4.1 Autoregressive receptive fields for noise
$p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{i=1}^{N}\prod_{j=1}^{M}p_{\theta}(x_{i,j}|\mathbf{z},x_{i,1},\dots,x_{i,j-1}).$ (5)
@@
4.2 Decoding the signal
Once our VAE has been trained, its latent space will represent clean signals, i.e., each $\mathbf{z}$ contains all the information about an $\mathbf{s}$. We would now like to use the model for denoising by inferring possible clean signals $\mathbf{s}$ for a given noisy image $\mathbf{x}$. Unfortunately, unlike previous methods [36, 35, 39], we cannot directly sample clean images from our encoder. Rather, samples from $q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})$ will directly correspond to a clean signal $\mathbf{s}$. We denote the signal corresponding to a value of $\mathbf{z}$ as $\mathbf{s}(\mathbf{z})$. For an experimental validation of this deterministic relationship, please refer to the supplementary material. To obtain denoised images, we approximate $\mathbf{s}(\mathbf{z})$ with an additional regression network, termed the signal decoder, $f_{\nu}(\mathbf{z})\approx\mathbf{s}(\mathbf{z})$. In the following, we describe how $f_{\nu}(\mathbf{z})$ is trained despite not having access to training pairs $(\mathbf{z}^{k},\mathbf{s}(\mathbf{z}^{k}))$.
In fact, training the signal decoder only requires pairs $(\mathbf{z}^{k},\mathbf{x}^{k})$, where $\mathbf{x}^{k}$ is a noisy image and $\mathbf{z}^{k}\sim q_{\phi}(\mathbf{z}|\mathbf{x}^{k})$. If we assume that the approximate posterior is accurate, such that $q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})\approx p_{\mathbf{\theta}}(\mathbf{z}|\mathbf{x})$, these pairs can be equivalently thought of as sampled from the joint distribution, $(\mathbf{z}^{k},\mathbf{x}^{k})\sim p_{\theta}(\mathbf{z},\mathbf{x})$.
Least squares regression analysis tells us that optimizing the signal decoder with the $L2$ loss,
$\mathcal{L}(\nu)=|f_{\nu}(\mathbf{z}^{k})-\mathbf{x}^{k}|^{2},$ (6)
will train it to approximate $\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{z})}[\mathbf{x}]$ for any $\mathbf{z}$ [13]. Since $\mathbf{z}$ contains no more or less information about $\mathbf{x}$ than $\mathbf{s}(\mathbf{z})$, the signal decoder will equivalently be approximating $\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{s}(\mathbf{z}))}[\mathbf{x}]$. Recalling that imaging noise is zero-centered (Sec. 2.1), the expected value of a noisy image given an underlying signal is that signal. Therefore, $f_{\nu}(\mathbf{z})\approx\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{s}(\mathbf{z}))}[\mathbf{x}]=\mathbf{s}(\mathbf{z})$. This is similar to Noise2Noise [27], where the regressor for $\mathbf{s}$ is trained using noisy targets $\mathbf{x}$.
@@ 4.2 Decoding the signal
4.3 Inference
With both the VAE and signal decoder trained, we can denoise an image $\mathbf{x}$ in a two step process. We first sample a latent variable $\mathbf{z}\sim q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})$ and then obtain a clean image by decoding $\hat{\mathbf{s}}=f_{\nu}(\mathbf{z})$. Similarly to [36, 35, 39] the result constitutes a random possible solution $\hat{\mathbf{s}}\sim p(\mathbf{s}|\mathbf{x})$. To obtain a consensus solution, we follow [36, 35, 39] in averaging a large number of such samples. In the following experiments, all results are the mean of 100 samples, both for our method and the baseline HDN36.
@@ 5 Experiments
Table 1: Comparison with baseline methods using Peak Signal-to-Noise Ratio (PSNR), higher is better. Best results for methods that do not require clean images are printed in bold, and best results overall are underlined. The dataset that contains spatially uncorrelated noise and the methods that address only unstructured noise are marked with an asterisk*. CARE is a supervised denoiser requiring examples of clean images, unlike the other baselines.
@@ 5.1 Datasets
We tested the performance of our proposed denoiser on real noisy images captured by four different imaging modalities that commonly suffer from row-correlated noise. The first is the EMCCD sensor for which we have three fluorescence microscopy datasets with known ground truth: Convallaria A [2], Convallaria B [37] and Mouse Actin [37]. The second is Scanning Confocal Microscopy (SCM), for which we have three fluorescence microscopy datasets with known ground truth: Mouse Nuclei [37], Actin Confocal [12] and Mito Confocal [12]. Note that Mouse Nuclei contains spatially uncorrelated noise, but was included to demonstrate that the proposed method is still applicable to spatially uncorrelated noise without modification. There is also the sCMOS sensor for which we have one fluorescence dataset with unknown ground truth: Embryo [9], followed by the microbolometer for which we have one infrared imaging dataset with unknown ground truth: IR [34]. Lastly, we include a STEM dataset with unknown ground truth: STEM [15]. For details on dataset size and train/test splits, please see each original publication.
@@
In addition, we created two datasets by corrupting the Flickr Faces HQ thumbnails dataset [18] with simulated noise. For FFHQ - Stripes, the images were corrupted by a combination of additive white Gaussian noise, Poisson shot noise and additive white Gaussian noise that had undergone a horizontal Gaussian blur. For FFHQ - Checkerboard, the images were corrupted by a combination of a vertical checkerboard pattern and Gaussian noise with an inverse signal dependence. For full details on how the simulated noise cases were created, please refer to the supplementary.
5.2 Baselines and architecture
@@
There is a scarcity in the literature of deep learning-based denoisers that can be applied to signal-dependent row-correlated noise and do not require examples of clean images or extensive calibration data. Currently, the only methods that meet these requirements are the self-supervised denoiser Structured Noise2Void (SN2V) [2] and the unsupervised denoiser Hierarchical DivNoising36 (HDN36[35]. We use these as baseline methods along with the supervised denoiser Content Aware Image Restoration (CARE) [44] for datasets where paired training images are available. For details on how each baseline is implemented in this paper, please refer to the supplementary.
@@
Of the denoisers not requiring paired images, HDN36’s performance is best, but the model implemented in the baseline’s publication uses significantly fewer parameters than ours (7 million to 25 million). We therefore evaluate an additional version, termed HDN36 large, with a similar number of parameters made by increasing the number of latent dimensions from 32 to 64.
@@
It should be noted that the Convallaria B and Mouse Actin datasets had been treated as spatially uncorrelated by Prakash et al. [35] when testing the denoiser HDN. We additionally report those results in Tab. 1. It should also be noted that the noise in the Mouse Nuclei dataset is spatially uncorrelated, so was denoised by HDN and its higher parameter version HDN large, which was made in the same way as HDN36 large.
@@
As for our model’s encoder, latent variables are produced by a fully convolutional hierarchical VAE [40], with 14 levels to its hierarchy. In addition to the full sized version, we evaluate a smaller version that requires approximately 6GB (as opposed to 20GB) of GPU memory to train. This was achieved by reducing the number of latent variables from 14 to 6 and reducing the number of latent dimensions from 64 to 32. We refer to these models as Ours small and Ours large, for the lower memory and higher memory versions respectively. Please see the supplementary materials for full architecture and training details.
5.3 Comparing denoising performance
Figure 3: Visual results from our method and the two unpaired baselines on all datasets. The spatial autocorrelation of the noise is overlaid on each noisy image, with red indicating positive correlation and blue indicating negative correlation. The direction of the correlation is given by the orientation of the autocorrelation bar. Additionally, the signal dependence of the noise in each dataset is shown in the graphs in the right-hand column. On the horizontal axis of these graphs is the clean signal intensity as a percentage of the maximum, while on the vertical axis is the variance of noisy pixel values recorded for these signal intensities. Ground truth must be used to calculate signal dependence, so orange lines are used where denoised images from our method are used as pseudo-ground truth.
Quantitative results measuring the Peak Signal-to-Noise Ratio (PSNR) in dB are reported in Tab. 1 and qualitative results are reported in Fig. 3. Note that HDN36 failed to train with the Mouse Actin and the Embryo dataset. Out of the methods that do not require paired images, Ours large achieved the highest PSNR across all datasets, even beating the supervised CARE on four of six microscopy datasets. Ours small then had the second highest PSNR for an unpaired method on all datasets except FFHQ - Stripe, where it was beaten by HDN36 large, although it should be noted that HDN36 large requires almost 20GB of GPU memory to train while Ours small requires only 6GB.
Turning to the qualitative results in Fig. 3, we see that Ours large denoised images from each dataset without leaving behind any artifacts, whereas HDN36 large could not remove the correlated component of the noise from the STEM or the FFHQ - Checkerboard dataset. SN2V left artifacts on all datasets with spatially correlated noise, but not for the uncorrelated Mouse Nuclei dataset. It can also be seen that Ours large produced the sharpest and most accurate images, with the exception of Mito Confocal, for which HDN36 large was sharper.
@@ 5.4 Ablation study - receptive field size
Figure 4: Here, we denoised the FFHQ - Checkerboard dataset 5 times, varying the number of pixels covered by the AR decoder’s receptive field. The PSNR of images denoised by each model was then calculated. Images show denoising results for different receptive field sizes with PSNR overlaid. Additionally, an image denoised using a full AR receptive field is included on the right along with its PSNR. In this situation, the signal decoder is given completely uninformative inputs and learns to output the mean of the entire training dataset.
As stated in Sec. 4, to model the noise correlations addressed in this paper, the receptive field of a VAE’s AR decoder must span pixels in the same row or column as the pixel being predicted. To effect of the investigate receptive field size (number of pixels), we denoised the FFHQ - Checkerboard dataset using a range of RF lengths, from 10 pixels to 120 pixels, and measured the effect on PSNR. We also include the effect of training a model with a full AR receptive field, as shown in Fig. 2b. The results are reported in Fig. 4.
The study shows that the AR decoder is able to model this spatially correlated noise, and therefore have the noise removed by the encoder, when the receptive field spans 40 pixels. Moreover, the decoder does not seem to model more of the signal as the receptive field grows. If it did, we would expect a steady drop in PSNR as denoised images lose signal content.
However, as the image on the right of Fig. 4 shows, the signal will be modeled by an AR decoder with a full receptive field. Latent variables will therefore carry no information about $\mathbf{x}$, causing the signal decoder to minimize Eq. 6 by predicting the mean of the training set. This is to be expected, as when the latent variable $\mathbf{z}$ carries no information about $\mathbf{x}$, $\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{z})}[\mathbf{x}]=\mathbb{E}_{p_{\theta}(\mathbf{x})}[\mathbf{x}]$.

### v3

@@ 4 Method
Figure 2: a): A Variational Autoencoder [5] (solid arrows) is trained to model the distribution of noisy images $\mathbf{x}$. The autoregressive (AR) decoder models the noise component of the images, while the latent variable models only the clean signal component $\mathbf{s}$. In a second step (dashed arrows), our novel signal decoder is trained to map latent variables into image space, producing an estimate of the signal underlying $\mathbf{x}$. b): To ensure that the decoder models only the imaging noise and the latent variables capture only the signal, we modify the AR decoder’s receptive field. In a full AR receptive field (Eq. 4), each output pixel (red) is a function of all input pixels located above and to the left (blue). In our decoder’s row-based AR receptive field (Eq. 5), each output pixel is a function of input pixels located in the same row, which corresponds to the row-correlated structure of imaging noise.
We propose a VAE-based unsupervised image denoiser for noise that is both signal-dependent and correlated along rows or columns of pixels. It is trained using only noisy images and does not require a pre-trained noise model. Moreover, it does not require pixel blind-spots or subsampling. Instead, we restrict the receptive field of the AR decoder in a hierarchical VAE [49] to a row or column of pixels, allowing the AR decoder to model the correlations of noise content but not the correlations of underlying signal. Following the insights of Chen et al. [5] (Sec. 3.3), the AR decoder will therefore learn to model noise, leaving the underlying signal to be encoded in the latent variable $\mathbf{z}$. We then propose a method for taking these latent variables and mapping them back into image space, obtaining estimates of clean signals. An outline of our method can be found in Fig. 2.
Autoregressive Receptive Fields for Noise Our AR decoder has a one-dimensional receptive field that is sufficient for modeling row- or column-correlated noise (Eq. 2) as it occurs in microscopy (see Sec. 3.1 Row-Correlated Noise) by spanning pixels in the same row or column as $x_{i,j}$. See Fig. 2b for a visual representation.
@@
$p_{\mathbf{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{i=1}^{N}\prod_{j=1}^{M}p_{\theta}(x_{i,j}|\mathbf{z},x_{i,1},\dots,x_{i,j-1}).\vspace{-1mm}$ (5)
@@
Decoding the Signal Once our VAE has been trained, its latent space will represent clean signals, i.e., each $\mathbf{z}$ contains all the information about an $\mathbf{s}$. We would now like to use the model for denoising by inferring possible clean signals $\mathbf{s}$ for a given noisy image $\mathbf{x}$. Unfortunately, unlike previous methods [45, 44, 48], we cannot directly sample clean images from our encoder. Rather, samples from $q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})$ will directly correspond to a clean signal $\mathbf{s}$. We denote the signal corresponding to a value of $\mathbf{z}$ as $\mathbf{s}(\mathbf{z})$. For an experimental validation of this deterministic relationship, please refer to the supplementary material. To obtain denoised images, we approximate $\mathbf{s}(\mathbf{z})$ with an additional regression network, termed signal decoder, $f_{\nu}(\mathbf{z})\approx\mathbf{s}(\mathbf{z})$. In the following, we describe how $f_{\nu}(\mathbf{z})$ is trained despite not having access to training pairs $(\mathbf{z},\mathbf{s}(\mathbf{z}))$.
In fact, training the signal decoder only requires pairs $(\mathbf{z}_{k},\mathbf{x}_{k})$, where $\mathbf{x}_{k}$ is a noisy image and $\mathbf{z}_{k}\sim q_{\phi}(\mathbf{z}|\mathbf{x}_{k})$. If we assume that the approximate posterior is accurate, such that $q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})\approx p_{\mathbf{\theta}}(\mathbf{z}|\mathbf{x})$, these pairs can be equivalently thought of as samples from the joint distribution, $(\mathbf{z}_{k},\mathbf{x}_{k})\sim p_{\theta}(\mathbf{z},\mathbf{x})$.
Least squares regression analysis tells us that optimizing the signal decoder by minimizing the squared $L2$ norm,
$\mathcal{L}(\nu)=\lVert f_{\nu}(\mathbf{z}_{k})-\mathbf{x}_{k}\rVert_{2}^{2},$ (6)
will train it to approximate $\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{z})}[\mathbf{x}]$ for any $\mathbf{z}$ [17]. By definition, $\mathbf{z}$ contains no more or less information about $\mathbf{x}$ than $\mathbf{s}(\mathbf{z})$, so the signal decoder will equivalently be approximating $\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{s}(\mathbf{z}))}[\mathbf{x}]$. Recalling that imaging noise is zero-centered (Sec. 3.1), the expected value of a noisy image given an underlying signal is that signal. Therefore, $f_{\nu}(\mathbf{z})\approx\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{s}(\mathbf{z}))}[\mathbf{x}]=\mathbf{s}(\mathbf{z})$. This is similar to Noise2Noise [33], where the regressor for $\mathbf{s}$ is trained using noisy targets $\mathbf{x}$.
@@
Inference  With the VAE and signal decoder trained, we can denoise an image $\mathbf{x}$ in a two-step process. We first sample a latent variable $\mathbf{z}\sim q_{\mathbf{\phi}}(\mathbf{z}|\mathbf{x})$, then obtain a clean image by decoding $\hat{\mathbf{s}}=f_{\nu}(\mathbf{z})$. Similarly to [45, 44, 48], the result constitutes a random possible solution $\hat{\mathbf{s}}\sim p(\mathbf{s}|\mathbf{x})$. To obtain a consensus solution, we follow [45, 44, 48] in averaging a large number of such samples. In the following experiments, all the results are the mean of 100 samples, both for our method and the baseline HDN36. Please refer to the supplementary material for the inference time and PSNR of the mean of 1, 10 and 1000 samples.
Figure 3: Visual results from our method and two unpaired baselines on all datasets. The spatial autocorrelation of the noise is overlaid on each noisy image, with red indicating a positive correlation and blue indicating a negative correlation. The direction of the correlation is given by the orientation of the autocorrelation bar. Additionally, the signal dependence of the noise in each dataset is shown in the right-hand column. The horizontal axis of these graphs is the clean signal intensity as a percentage of the maximum, while the vertical axis is the variance of noisy pixel values recorded for these signal intensities. Ground truth must be used to calculate signal dependence, so orange lines are used where denoised images from our method are used as pseudo-ground truth.
@@ 5 Experiments
Table 1: Comparison to baseline methods using Peak Signal-to-Noise Ratio (PSNR), higher is better. The best results for methods that do not require paired images are printed in bold, and best results overall are underlined. HDN is a variant of HDN36 that is intended for spatially uncorrelated noise, but was applied to two datasets with spatially correlated noise by the original authors. These extra results are presented in brackets. Note that the Mouse Nuclei dataset contains spatially uncorrelated noise and that HDN36 failed to train on the Mouse Actin dataset. CARE and N2N are supervised denoisers requiring paired images, unlike the other baselines.
@@ 5.1 Benchmarking Denoising Performance
Datasets We tested the performance of our proposed denoiser on real noisy images captured by five different imaging modalities that commonly suffer from row-correlated noise. Two modalities have noisy image datasets with known ground truth for quantitative evaluation. The method for obtaining ground truth can be found in the source publications. The first is the EMCCD sensor, for which we have three spinning disk confocal fluorescence microscopy datasets: Convallaria A [3], Convallaria B [46] and Mouse Actin [46]. The second is LSCM, for which we have three fluorescence datasets: Mouse Nuclei [46], Actin Confocal [16] and Mito Confocal [16]. Note that Mouse Nuclei contains spatially uncorrelated noise, but was included to demonstrate that the proposed method is still applicable to spatially uncorrelated noise without modification. For details on the size of the datasets and the train/test splits, see original publications. The remaining modalities have only noisy images. These are the sCMOS sensor for which we have one fluorescence dataset: Embryo [13], the microbolometer with one infrared imaging dataset: IR [43], and a STEM dataset: STEM [19].
@@
In addition to real data, we created two datasets by corrupting the Flickr Faces HQ thumbnails dataset [24] with simulated noise. For FFHQ - Stripes, the images were corrupted by a combination of additive white Gaussian noise, Poisson shot noise, and additive white Gaussian noise that had undergone a horizontal Gaussian blur. For FFHQ - Checkerboard, the images were corrupted by a combination of a vertical checkerboard pattern and Gaussian noise with an inverse signal dependence. Details on the simulated noises are in the supplementary materials.
@@
Baselines and Architecture We compare our method with other deep learning-based denoisers that can be applied to spatially correlated, signal-dependent noise and do not require paired images. These are the self-supervised denoisers Structured Noise2Void (SN2V) [3], Noise2Void2 (N2V2) [21] and Asymmetric PD BSN (AP-BSN) [32], and the unsupervised denoiser Hierarchical DivNoising36 (HDN36[44]. We use these as baseline methods along with Content-Aware Image Restoration (CARE) [56], which is trained using pairs of noisy and clean images, and Noise2Noise (N2N) [33], which is trained using two noisy acquisitions of the same signal, on datasets with suitable training images available. For details on how each baseline was implemented, refer to the supplementary material.
@@
Of the denoisers not requiring paired images, the performance of HDN36 is the best, but the model implemented in the baseline’s publication uses significantly fewer parameters than ours (7 million to 25 million). We therefore evaluate an additional version, termed HDN36 large, with a similar number of parameters made by increasing the number of latent dimensions from 32 to 64.
@@
It should be noted that the Convallaria B and Mouse Actin datasets had been treated as spatially uncorrelated by Prakash et al. [44] when HDN was tested. We also report those results in Tab. 1. It should also be noted that the noise in the Mouse Nuclei dataset is spatially uncorrelated, so was denoised by HDN and its higher parameter version, HDN large, which was made the same way as HDN36 large.
@@
As for our model’s encoder, latent variables are produced by a hierarchical VAE [49] with 14 levels to its hierarchy. In addition to the full-sized version, we evaluate a smaller version that requires approximately 6GB (as opposed to 20GB) of GPU memory to train. This was achieved by reducing the number of latent variables from 14 to 6 and reducing the number of latent dimensions from 64 to 32. We refer to these models as Ours small and Ours large, for the lower memory and higher memory versions respectively. See the supplementary materials for full architecture and training details.
Discussion Quantitative results measuring the Peak Signal-to-Noise Ratio (PSNR) in dB are reported in Tab. 1, and qualitative results are reported in Fig. 3. Note that HDN36 failed to train with the Mouse Actin and the Embryo dataset. Out of the methods that do not require paired images, Ours large achieved the highest PSNR across all datasets, even beating the supervised CARE and N2N on four of six microscopy datasets. Ours small then had the second highest PSNR for an unpaired method on all datasets except FFHQ - Stripe .
Turning to the qualitative results in Fig. 3, we see that Ours large denoised images from each dataset without leaving behind any artifacts, whereas HDN36 large could not remove the correlated component of the noise from the STEM or the FFHQ - Checkerboard dataset. SN2V left artifacts in Convallaria A and FFHQ - Checkerboard. It is unclear if the high frequency features that are visible in the Mito Confocal output of HDN36 large and N2V2, but not in our output, are true signal or artifacts of remaining noise, as the ground truth also contains a low level of noise.
@@ 5.2 Ablation Study - Receptive Field Size
Figure 4: The FFHQ - Checkerboard dataset wass denoised 5 times by varying the number of pixels covered by the AR decoder’s receptive field. The images show denoising results for each receptive field size with PSNR overlaid. Additionally, an image denoised using a full AR receptive field is included on the right. In this situation, the signal decoder is given completely uninformative inputs and learns to output the mean of the training dataset.
As stated in Sec. 4, to model the noise correlations addressed in this paper, the receptive field of a VAE’s AR decoder must span pixels in the same row or column as the pixel being predicted. In Fig. 4, the effect of receptive field size (number of pixels) is investigated by denoising the FFHQ - Checkerboard dataset using a range of receptive field lengths, from 10 to 120 pixels, and measuring the effect on PSNR. We also include the effect of training a model with a full AR receptive field, as shown in Fig. 2b.
This study shows that the AR decoder is able to model this spatially correlated noise, and therefore have it removed by the encoder, when the receptive field spans 40 pixels. Moreover, the AR decoder does not seem to model more of the signal as the receptive field grows. If it did, we would expect a steady drop in PSNR as denoised images lose signal content. However, as the image on the right of Fig. 4 shows, the signal will be modeled by an AR decoder with a full receptive field. Latent variables in this situation carry no information about $\mathbf{x}$, causing the signal decoder to minimize Eq. 6 by predicting the mean of the training set. This is expected, as when the latent variable $\mathbf{z}$ has no information about $\mathbf{x}$, $\mathbb{E}_{p_{\theta}(\mathbf{x}|\mathbf{z})}[\mathbf{x}]=\mathbb{E}_{p_{\theta}(\mathbf{x})}[\mathbf{x}]$.